# SynSUM – Synthetic Benchmark with Structured and Unstructured Medical Records

### v2

@@ 5 Discussion
The dataset is primarily designed to facilitate research on clinical information extraction in the presence of tabular background variables. Future work will focus on realizing the idea presented in Figure 2, where the tabular features aid in more accurately extracting concepts from the text by linking them through domain knowledge. Apart from this, we also foresee multiple secondary uses of the dataset. First, the dataset could facilitate research on the automation of clinical reasoning over tabular data and text, following the example of Rabaey et al. (2024). Second, it could be used to benchmark causal effect estimation methods in the presence of textual confounders, similar to Veitch, Sridhar, and Blei (2020), thanks to the purposeful inclusion of both a treatment and outcome variable in our dataset. Third, there has been increasing interest in clinical synthetic data (Hernandez et al. 2022), where a set of patient characteristics is turned into a synthetic version that is meant to protect the privacy of individuals in the original dataset. Our dataset could serve as a benchmark for comparing synthetic data generation methods that jointly generate tabular variables and text (Lee 2018; Ceritli et al. 2023; Guan et al. 2019). In short, any area of research focusing on the intersection of tabular data and text in healthcare can potentially benefit from our proposed benchmark.
Limitations
While the dataset we constructed is meant to have realistic properties, we also intentionally simplify reality to make the design and generation process feasible. The dataset is purely meant as a research benchmark where the ground truth relations are known, and results obtained on it are not meant to transfer to real clinical notes or datasets. This is confirmed by the evaluators in our expert study, who warned that the writing style used in the text notes is not in accordance with reality, even if the content is mostly realistic. We therefore advise strongly against using the dataset for training prediction models which will be deployed in real settings. Furthermore, our dataset contains static patients that do not evolve over time. While this is a reasonable simplification to make in the realm of primary care, time series are a crucial modality in more complex settings, such as intensive care.
Acknowledgments
@@
Paloma Rabaey and Henri Arno’s research is funded by the Research Foundation Flanders (FWO Vlaanderen) with grant numbers 1170124N and 11Q2C24N. This research also received funding from the Flemish government under the “Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen” programme. The authors would like to thank Géraldine Deberdt, Thibault Detremerie, An De Sutter, Veerle Piessens en Florian Stul for participating in the expert evaluation of the artificial clinical notes.
Appendix
Appendix A Clinical Information Extraction with Background Knowledge
Clinical information extraction can be improved by exploiting tabular background information that is present in the EHR, next to the text from which we are extracting clinical concepts (e.g. the symptoms). Figure 2 illustrates this idea with two practical examples. The SynSUM benchmark enables future research to explore this idea, by linking structured tabular features with unstructured text describing a fictional patient encounter.
Appendix B Bayesian network
$\mathcal{P}_{\text{{joint}}}(\text{{asthma, smoking, ..., antibio, \# days}})=\mathcal{P}(\text{{asthma}})\mathcal{P}(\text{{smoking}})\mathcal{P}(\text{{COPD}}\mid\text{{smoking}})\\ \mathcal{P}(\text{{hayf}})\mathcal{P}(\text{{season}})\mathcal{P}(\text{{pneu}}\mid\text{{asthma, COPD, season}})\mathcal{P}(\text{{cold}}\mid\text{{season}})\\ \mathcal{P}(\text{{dysp}}\mid\text{{asthma, smoking, COPD, pneu, hayf}})\mathcal{P}(\text{{cough}}\mid\text{{asthma, smoking, COPD, pneu, cold}})\\ \mathcal{P}(\text{{pain}}\mid\text{{cough, pneu, COPD, cold}})\mathcal{P}(\text{{fever}}\mid\text{{pneu, cold}})\\ \mathcal{P}(\text{{nasal}}\mid\text{{cold, hayf}})\mathcal{P}(\text{{policy}})\mathcal{P}(\text{{self-empl}})\mathcal{P}(\text{{antibio}}\mid\text{{policy, dysp, cough, pain, fever}})\\ \mathcal{P}(\text{{\#days}}\mid\text{{antibio, dysp, cough, pain, fever, nasal, self-empl}})$ (1)
Name Type Description Values
Asthma underlying condition Chronic lung disease in which the airways narrow and swell yes/no
Smoking underlying condition Whether the patient is a regular smoker of tobacco yes/no
COPD underlying condition Chronic Obstructive Pulmonary Disease, where airflow from the lungs is obstructed yes/no
Hay fever underlying condition Allergic rhinitis, irritation of the nose caused by an allergen (e.g. pollen) yes/no
Season non-clinical Season of the year winter/summer
Pneumonia diagnosis Infection that inflames the air sacs in one or both lungs yes/no
Common cold diagnosis Upper respiratory tract infection, irritation and swelling of the upper airways yes/no
Dyspnea symptom Shortness of breath, the feeling of not getting enough air yes/no
Cough symptom Any type of cough, no distinction between non-productive (dry) or productive (bringing up mucus or phlegm) yes/no
Pain symptom Pain related to the airways or chest area yes/no
Fever symptom Elevation of body temperature high/low/none
Nasal symptom Nasal symptoms, such as runny nose or sneezing yes/no
Antibiotics treatment Whether any type of antibiotics are prescribed to the patient yes/no
# Days at home outcome How many days the patient ends up staying home as a result of their symptoms and treatment discrete $(0-...)$
Table 3: Description of tabular variables in our dataset.
To define a data generating mechanism from which we can sample synthetic patients, we turn the DAG from Figure 1 into a Bayesian network by defining a joint probability distribution. In a Bayesian network, this joint distribution factorizes into the product of conditional probability distributions for each variable, as shown in Equation (1). We parameterize these using four different approaches.
Conditional probability table
When the variable is discrete and has a limited number of parents, we define a conditional probability table (CPT). Each entry of the table contains the probability for a particular value of the variable, conditional on the combination of values of the parent variables. If the variable has no parents, we just define a prior probability. The probabilities in the tables were filled in by the expert based on experience, as well as demographics in Belgium and the expert’s local general practice. While we do not expect these probabilities to generalize to the global patient population as a whole, a realistic-looking distribution suffices for our use-case. We provide these tables for the variables asthma, smoking, hay fever, COPD, season, pneumonia, common cold, fever, policy and self-employed in Figure 3.
Noisy-OR distribution
For categorical variables with many parents, it becomes infeasible to manually fill in the CPT in a clinically meaningful way, because of the large number of possible combinations of parent values. This is the case for the symptoms dyspnea, cough, pain and nasal in our Bayesian network. To circumvent this problem, we define a Noisy-OR distribution (Koller and Friedman 2009). The Noisy-OR model is commonly used to define the distribution of a variable $Y$ which depends on a set of causes $\{X_{1},\dots,X_{k}\}$. It rests on the assumption that the combined influence of the possible causes $\{X_{1},\dots,X_{k}\}$ on $Y$ is a simple combination of the influence of each $X_{i}$ on $Y$ in isolation. This is a reasonable assumption to make in the case of symptoms with multiple possible causes (parents in the Bayesian network): a symptom arises in a patient if any of its possible causes succeeds in activating the symptom through its own independent mechanism. As shown in Equation (2), the parameterization of the noisy-OR distribution rests on choosing the parameters $p_{i}$, which is the probability that a possible cause $X_{i}$ activates symptom $Y$. As a special case, $p_{0}$, also known as the leak probability, is the probability that symptom $Y$ is activated as the result of another unmodeled cause (ouside of all $X_{i}$’s). Note that $x_{i}$ in the equation is $1$ when the cause $X_{i}$ is present in the patient, and 00 if not. Equations (3) through (6) define such a Noisy-OR distribution for the symptoms dyspnea, cough, pain and nasal. Note that the symptom fever is fully defined through a CPT, since the expert was able to provide intuition on all possible combinations of its two parent values, eliminating the need for a Noisy-OR distribution.
$\operatorname{Noisy-OR}(p_{0},p_{1},\dots,p_{k})\coloneq\\ \mathcal{P}(Y=1\mid X_{1},\dots,X_{k})\\ =1-\mathcal{P}(Y=0\mid X_{1},\dots.,X_{k})\text{, with}\\ \mathcal{P}(Y=0\mid X_{1}=x_{1},\dots,X_{k}=x_{k})\\ =(1-p_{0})(1-p_{1})^{x_{1}}\dots(1-p_{k})^{x_{k}}\\ \text{and }x_{1}...x_{k}\in\{0,1\}$ (2)
$\mathcal{P}(\text{{dysp}}\mid\text{{asthma, smoking, COPD, hayf, pneu}})\\ =\operatorname{Noisy-OR}(p_{0}=0.05,p_{\text{{asthma}}}=0.9,p_{\text{{smoking}}}=0.3,\\ p_{\text{{COPD}}}=0.9,p_{\text{{hayf}}}=0.2,p_{\text{{pneu}}}=0.3)$ (3)
$\mathcal{P}(\text{{cough}}\mid\text{{asthma, smoking, COPD, pneu, cold}})\\ =\operatorname{Noisy-OR}(p_{0}=0.07,p_{\text{{asthma}}}=0.3,p_{\text{{smoking}}}=0.6,\\ p_{\text{{COPD}}}=0.4,p_{\text{{pneu}}}=0.85,p_{\text{{cold}}}=0.7)$ (4)
$\mathcal{P}(\text{{pain}}\mid\text{{COPD}},\text{{cough}},\text{{pneu}},\text{{cold}})\\ =\operatorname{Noisy-OR}(p_{0}=0.05,p_{\text{{COPD}}}=0.15,p_{\text{{cough}}}=0.2,\\ p_{\text{{pneu}}}=0.3,p_{\text{{cold}}}=0.1)$ (5)
$\mathcal{P}(\text{{nasal}}\mid\text{{hayf, cold}})\\ =NoisyOR(p_{0}=0.1,p_{\text{{hayf}}}=0.85,p_{\text{{cold}}}=0.7)$ (6)
Logistic regression
Whether or not to prescribe antibiotics depends on whether the clinician suspects pneumonia in the patient. Their suspicion raises with the number of symptoms present in the patient, with some symptoms weighing more than others. Once their level of suspicion reaches a certain threshold, they decide to prescribe treatment. This process can be modeled using a logistic regression model taking the symptoms dyspnea, cough, pain and fever, as well as the variable policy, as an input, as shown in Equation (7). Here, $x_{po}$ (policy) can take on the values 1 (high) or 0 (low), $x_{d}$ (dyspnea), $x_{c}$ (cough) and $x_{pa}$ (pain) can take on the value 1 (yes) or 0 (no), and $x_{f}$ (fever) can be 2 (high), 1 (low) or 0 (none). The bias of -$\text{-}3$ was set based on the following constraint: if there’s no symptoms at all, and policy is low, then the probability of prescribing antibiotics (due to some other unmodeled cause) should be around 5%. Similarly, the coefficient for policy was set to fit the following constraint: if there’s no symptoms at all, and policy is high, the probability should be around 10%. All other coefficients were then chosen by the expert based on the relative importance of the symptoms when deciding to prescribe antibiotics, taking the coefficient for policy as a starting point. As a final sanity-check, we asked the expert to label a set of test cases with whether they would prescribe antibiotics or not, allowing us to compare with the probability predicted by the model. Table 4 shows these results. We see that the predictions made by the model mostly correspond well with the clinician’s intuition, confirming that the proposed coefficients make sense.
$\mathcal{P}(\text{{antibio}}=\text{{yes}}\mid\text{{policy}}=x_{po},\text{{dysp}}=x_{d},\\ \text{{cough}}=x_{c},\text{{pain}}=x_{pa},\text{{fever}}=x_{f})\\ =\text{Sigmoid}(-3+1\times x_{po}+0.8\times x_{d}+0.665\times x_{c}\\ +0.665\times x_{pa}+0.9\times(x_{f}==1)+2.25\times(x_{f}==2)),\\ \text{with }x_{po},x_{d},x_{c},x_{pa}\in\{0,1\}\text{, and }x_{f}\in\{0,1,2\}$ (7)
@@ Appendix B Bayesian network
[…]
Poisson regression
Finally, we need to model the number of days the patient ends up staying home due to their complaints. This depends on the symptoms experienced by the patient, as well as whether they received antibiotics as a treatment. Since the outcome is discrete, with most patients staying home for a low number of days, we decided to model this using a Poisson regression. Assuming that the effect of getting treatment would be non-linear in relation to the presence or absence of the symptoms, we defined two separate Poisson models: one where no antibiotics were prescribed (Equation (8)), and one where they were prescribed (Equation (9)). Both models take the symptoms dysp, cough, pain, nasal and fever as an input, as well as the variable self-employed, and predict a mean number of days $\lambda$, which parameterizes the Poisson distribution. The coefficients for each model were tuned using gradient descent based on the train cases shown in Table 5. Like before, the expert was asked to (loosely) label these cases for how long they suspected the patient to stay home on average as a result of these symptoms. The coefficient for the variable self-employed was tuned manually, based on the assumption that being self-employed would shave some days off the predicted number, regardless of the particular symptoms experienced by the patient. As a sanity check, we compared the mean number of days predicted by the model (parameter $\lambda$ in the Poisson model) with the number of days estimated by the expert for a small test set of cases which were not seen during training. The results are shown in Table 6.
$\mathcal{P}(\text{{\# days}}\mid\text{{dysp}}=x_{d},\text{{cough}}=x_{c},\text{{pain}}=x_{pa},\\ \text{{nasal}}=x_{n},\text{{fever}}=x_{f},\text{{self-empl}}=x_{se},\text{{antibio}}=\text{{no}})\\ =Poisson(\lambda_{0})\\ \lambda_{0}=exp(0.010+0.64\times x_{d}+0.35\times x_{c}+0.47\times x_{pa}\\ +0.011\times x_{n}+0.81\times(x_{f}==1)\\ +1.23\times(x_{f}==2)-0.5\times x_{se})\\ \text{with }x_{d},x_{c},x_{pa},x_{n},x_{se}\in\{0,1\}\text{, and }x_{f}\in\{0,1,2\}$ (8)
$\mathcal{P}(\text{{\# days}}\mid\text{{dysp}}=x_{d},\text{{cough}}=x_{c},\text{{pain}}=x_{pa},\\ \text{{nasal}}=x_{n},\text{{fever}}=x_{f},\text{{selfempl}}=x_{se},\text{{antibio}}=\text{{yes}})\\ =Poisson(\lambda_{1})\\ \lambda_{1}=exp(0.16+0.51\times x_{d}+0.42\times x_{c}+0.26\times x_{pa}\\ +0.0051\times x_{n}+0.24\times(x_{f}==1)\\ +0.57\times(x_{f}==2)-0.5\times x_{se})\\ \text{with }x_{d},x_{c},x_{pa},x_{n},x_{se}\in\{0,1\}\text{, and }x_{f}\in\{0,1,2\}$ (9)
@@
[…]
@@
[…]
Sampling
The joint probability distribution from Equation (1) is now fully specified. We can use this Bayesian network to randomly sample the tabular portion of a patient record top-down, starting from the root variables without parents at the top and continuing further down. Each value is sampled conditionally on the variable’s parents’ values, using the conditional distributions we have defined. We repeat this process $10{,}000$ times, leaving us with $10{,}000$ artificial patient records consisting of 16 tabular features.
Appendix C Prompting the large language model
Starting from the tabular portion of the patient record, we aim to generate a text describing this fictional patient encounter. The scenario we simulate artificially is as follows. The patient goes to the primary care physician, telling them their symptoms and possible underlying conditions, along with additional context on the severity of these symptoms, when they started, among other details. The physician takes descriptive notes during this consultation, writing down the (recent) history prescribed by the patient. Then, based on the patient’s described complaints, they conduct a physical examination, writing down all findings. Both parts together then form the textual description of the patient encounter.
Presence of symptoms
The first block of information in the prompt concerns the symptoms experienced by the patient. We do not list the full set of symptoms exhaustively. Even if a patient might experience a certain symptom, there is a possibility that they do not mention it to the clinician, or that the clinician does not find it noteworthy to write down. On the other hand, if a patient does not experience a symptom, it is not very likely that they will mention this to the physician, and the physician does not always have a reason to ask for the symptom either. We therefore ask our expert to list the probability of mentioning the symptom in a clinical note when the symptom is positive and when it is negative. Of course, this would not generalize to all physicians, but it helps to bring some variety and realism in the notes we generate. The probabilities are as follows:
For each symptom, we sample whether it is to be mentioned in the prompt, conditional on its value, according to the probabilities stated above. As can be seen in Figure 1, we explicitly tell the model what symptoms to mention and which to steer clear from. We randomly permute the ordering of the symptoms in each prompt.
Symptom descriptors
Symptom Cause Descriptors
smoking during exercise, worse in morning, mild
hay fever light, mild, stuffy feeling, all closed up
pneumonia light, mild, severe, no clear cause
cough asthma attack-related, dry
smoking productive, mostly in morning, during exercise, gurgling
COPD phlegm, sputum, gurgling, worse when lying down
pain asthma tension behind sternum
COPD light, mild
Table 7: Descriptors used in the prompt to describe each symptom when it is present in the patient. Depending on the cause(s) of the symptom (as listed in the tabular patient record), we randomly sample from a different set of descriptors.
To make the note realistic, the LLM must invent some context regarding the patient’s symptoms when writing the history portion of the note. We want this context to indirectly relate to the cause of these symptoms, as they would in a real patient encounter. For example, a cough induced by asthma would likely be momentarily and attack-related, while a cough resulting from pneumonia might be more persistent over the longer term. We therefore ask the expert to write down a list of adjectives or phrases describing each symptom, conditioned on the cause of the symptom. These descriptors can be found in Table 7. The list of possible causes for a symptom is simply the list of parents in the Bayesian network.
For each symptom which is present in the patient and selected to be mentioned in the note, we check the tabular patient record for the possible causes. For example, for the symptom cough, the possible causes are asthma, smoking, COPD, pneumonia and common cold. In the example in Figure 1, asthma is the only cause which is “on”. We therefore randomly sample a descriptor from the list of descriptors for cough in the presence of asthma, in this case the adjective “dry”. This adjective is added in the prompt. If multiple causes are “on”, we find the strongest cause, and sample from that list. The strongest cause is pneumonia, followed by common cold, followed by all other causes. If neither pneumonia nor common cold is part of the multiple causes, we simply make a bag of all descriptors associated to the causes which are “on”, and sample from that bag. In the rare event that no causes are “on”, yet a symptom is still observed (which is possible due to the leak probability in the Noisy-OR distribution), we do not add a descriptor. Note that while the diagnoses pneumonia and common cold should not be mentioned explicitly, they indirectly and subtly influence the content of the note through the descriptors, adding another realistic dimension to the content of the note.
Underlying health conditions
While the diagnoses should not be mentioned directly in the note, it is realistic to assume that the note would mention underlying health conditions the patient may have. Since these health conditions are assumed to be known up-front, as they are part of the history of the patient, they may contribute a lot to the interpretation of the symptoms by both the patient themselves and the clinician writing down the note. We therefore add them to the prompt as well, as can be seen in Figure 1. We do not force the LLM to explicitly mention these in the note, since it seems feasible that a clinician would not mention them every time. Should there be more than one underlying condition, we mention them all, randomly permuting the order in each prompt. If there are no underlying health conditions, we simply remove this part of the prompt.
Additional instructions
We tell the LLM that the note must be structured with a “History” portion and a “Physical examination” portion. While the “History” portion describes the patient’s self-reported symptoms and underlying health conditions, which are in large part dictated by the prompt, the “Physical examination” portion leaves the LLM with more freedom to imagine additional clinical examinations which were performed on the patient. As such, the “Physical examination” portion has a lot of potential for adding complexity, clinical terminology and realism to the note.
We also add some additional instructions to the prompt, asking it not to mention any suspicions of possible diagnoses. We further tell the LLM it can imagine context or details, but no additional symptoms. We noticed that if we left this part out, the LLM would sometimes mention the symptoms we specifically asked to leave out. We ask not to mention patient gender or age, because preliminary testing revealed that the LLM often used the same age and gender (34-year old woman), which could confound or bias the notes. Finally, we add that the notes may be long (around 5 lines or more), to avoid the LLM being too succinct.
Special case: no respiratory symptoms
@@ Appendix C Prompting the large language model
Figure 4: Example of what would happen if we simply extended the general prompt to the 3629 cases where the patient does not experience any respiratory symptoms. There would be little variation in these notes, as these patients seem to visit the doctor’s office for no reason. It is more realistic to encourage the LLM to generate a note which describes a patient visiting for a non-respiratory complaint, like the prompts shown in Figures 5 and 6.
@@
There are 3629 out of 10,000 patients where all symptoms in the tabular record are “no”, meaning the patient does not experience any respiratory symptoms. If we used the same prompt as before, this would result in an unrealistic clinical note, since the note would simply list all symptoms the patient does not have, without giving an actual reason for the patient’s visit. Furthermore, there would be little variation in these notes. An example is shown in Figure 4. For these cases, it makes more sense to assume that the patient visits for a complaint unrelated to the respiratory domain, such as back pain, stomach issues, a skin rash, etc. To generate these special cases, we use a special prompt, telling the LLM the patient does not experience any of the 5 respiratory symptoms. When the patient has at least one underlying health condition (which is the case in 239 out of 3629 special cases), we add this to the prompt in the same way as before, like the example in Figure 5. If not (i.e., for the remaining 3390 out of 3629 special cases), we tell the LLM not to mention any of those health conditions either, see the prompt in Figure 6. The latter prompt asks for three clinical notes at once, encouraging the LLM to be more creative and not repeat the same scenario every time, as well as being a little more cost-effective. This is possible because of the prompt being non-specific to any of these 3390 patients. We then randomly distribute all generated texts to each tabular patient record within this subset.
Prompting details
As a large language model, we opted for OpenAI’s GPT-4o model, using the version released in May 2024 (OpenAI 2024). We set the temperature to 1.2 to encourage some more variation in the notes, while at the same time keeping them realistic. Before providing the case-specific prompt, we set the following system message: “You are a general practitioner, and need to summarize the patient encounter in a clinical note. Your notes are detailed and extensive.” We set the max_tokens parameter to 1000. All other parameters were kept as their default value. Generating all 10,000 notes and their compact version cost around $130$\$.
Appendix D Additional example prompts
Figures 7 and 8 show two additional example prompts.

### v3

@@ 3.3 Intended uses
Apart from this, we also foresee multiple secondary uses of the dataset. First, the dataset could facilitate research on the automation of clinical reasoning over tabular data and text, following the example of Rabaey et al. [24]. Second, it could be used to benchmark causal effect estimation methods in the presence of textual confounders, as has been done by Arno et al. [2]. This is possible thanks to the purposeful inclusion of both a treatment and outcome variable in our dataset. Third, there has been increasing interest in clinical synthetic data [7], where a set of patient characteristics is turned into a synthetic version that is meant to protect the privacy of individuals in the original dataset. Our dataset could serve as a benchmark for comparing synthetic data generation methods that jointly generate tabular variables and text [15, 5]. In short, any area of research focusing on the intersection of tabular data and text in healthcare can potentially benefit from our proposed benchmark.
Conversely, there are important limitations to the intended use of SimSUM. As it is fully synthetic—constructed from expert-defined tabular data and language model–generated notes—it does not capture the full complexity, variability, or noise characteristic of real electronic health records. For this reason, we strongly advise against using SimSUM to evaluate clinical model performance or to train models intended for production. Instead, its value lies in supporting research scenarios, such as those outlined above, which benefit from the availability of a known and controllable ground-truth structure.
4 Conclusions
We introduced SimSUM, a simulated dataset of structured and unstructured medical records designed to support research on clinical information extraction using background tabular data. By explicitly linking structured features and textual concepts through a Bayesian network, SimSUM provides a controlled setting to explore how domain knowledge can enhance information extraction from clinical notes. Indeed, our baseline results suggest that integrating tabular background features with text improves symptom extraction, particularly for harder-to-predict cases like pain and fever. Future research can explore more advanced hybrid models that leverage domain knowledge to link structured and unstructured data. Additionally, the annotated symptom spans, though imperfect, offer a foundation for developing lightweight span detection models and explainable extraction methods.
Limitations  Despite aiming for realism, SimSUM intentionally simplifies many aspects of real-world clinical data to support its role as a controlled research benchmark. The dataset is synthetic by design, with known ground-truth relationships, and is not intended to generalize to real clinical notes or settings. Expert reviewers confirmed that while the content of the notes is generally plausible, the writing style does not reflect how clinicians typically document patient encounters. Additionally, SimSUM represents static patient snapshots rather than evolving time series, which limits its applicability in domains like intensive care where temporal dynamics are critical.
Declarations
Funding
@@ 4 Conclusions
Paloma Rabaey’s research is funded by the Research Foundation Flanders (FWO Vlaanderen) with grant number 1170124N. This research also received funding from the Flemish government under the “Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen” programme.
Acknowledgements
The authors would like to thank Géraldine Deberdt, Thibault Detremerie, An De Sutter, Veerle Piessens and Florian Stul for participating in the expert evaluation of the artificial clinical notes. The authors would also like to thank Henri Arno for his early contributions to the dataset design.
Appendix A Bayesian network: Additional material
Table 7 shows the comparison of the probability predicted by the antibiotics model (Equation (7) in Section 2.1) with test cases labeled by the expert. Table 8 shows the expert-labeled train cases that were used to fit the coefficients in the # days model (Equation (8) and (9) in Section 2.1). Table 9 shows the comparison of the probability predicted by the # days model with test cases labeled by the expert.
[…]
@@ Appendix A Bayesian network: Additional material
Table 8: Train cases labeled by the expert on how many days they expect a patient to stay home (all assuming $\text{{self{\text{-}}employed}}=\text{{no}}$), with and without prescribing antibiotics. “label” indicates the expert’s estimation, while “pred” indicates the mean number of days $\lambda$ predicted by the #days models (Equation (8) and (9) in Section 2.1).
@@
Table 9: Test cases (not part of training set) labeled by the expert on how many days they expect a patient to stay home (all assuming $\text{{self{\text{-}}employed}}=\text{{no}}$), with and without prescribing antibiotics. “label” indicates the expert’s estimation, while “pred” indicates the mean number of days $\lambda$ predicted by the #days models (Equation (8) and (9) in Section 2.1).
@@
Appendix B Prompting strategy for special cases
We use an alternative prompting strategy to generate clinical notes for the patient records where none of the respiratory symptoms are present. This is the case for around one third of our dataset. In these cases, we encourage the LLM to imagine an alternative reason for the patient visit, unrelated to the respiratory domain. Fig. 5 shows a bad example of what would happen if we simply used the original prompt for these cases. Fig. 6 and Fig. 7 show the prompts we used instead, both for the cases where the patients had underlying conditions and where they did not.
@@ Appendix B Prompting strategy for special cases
Figure 5: Example of what would happen if we simply extended the general prompt to the 3629 cases where the patient does not experience any respiratory symptoms. There would be little variation in these notes, as these patients seem to visit the doctor’s office for no reason.
@@
Appendix C Additional example prompts
Fig. 8 and Fig. 9 show two additional example prompts.